OpenStudy (anonymous):

A system of equations is shown below: 8x + 5y = 9 3x + 2y = 4 Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

I have been on this forever please help! :(

OpenStudy (triciaal):

there are 4 ways to solve simultaneous equations here they told you which method to use replacing one equation with the sum of that equation and a multiple of the other the idea behind this is to cancel out (eliminate) one of the variables. you will be left with one equation with one variable. solve for that variable then substitute that value in the original to get the value of the other variable.

OpenStudy (anonymous):

I'm confused... :(

OpenStudy (triciaal):

8x + 5y = 9 equation 1 3x + 2y = 4 equation 2 let's multiply equation 1 by 2 and equation 2 by -5 post what you get we can label equation 3 and 4 respectively

OpenStudy (anonymous):

16x + 10y = 18

OpenStudy (triciaal):

good equation 3

OpenStudy (triciaal):

where is equation 4

OpenStudy (anonymous):

-15x + -10y = -20

OpenStudy (triciaal):

now add eqn 3 and eqn 4

OpenStudy (anonymous):

1x + 0y = -2

OpenStudy (triciaal):